OpenStudy (anonymous):

3x-y=20 x=7-y Solve for x and y.

OpenStudy (anonymous):

ill help :)

OpenStudy (mr.math):

\[3x-y=20\]+ \[x+y=7\] \[4x=27 \implies \dots\]

OpenStudy (mr.math):

Find \(x\) from this last equation, and then substitute back to find y.

OpenStudy (anonymous):

its really simple, all u gotta do is solve the equations throught the similultaneous method. there are 2 ways, 1. Substitution 2. Elimination

OpenStudy (anonymous):

They're both separate. It's not adding or subtracting them together...

OpenStudy (anonymous):

y=1/4;x=81/4

OpenStudy (anonymous):